Calculate the escape velocity from a red giant's atmosphere. Assume that the stars mass is 1 M and its radius is 100 R. How does this compare with the speed at which a planetary nebula shell is ejected?
I already calculate the escape velocity and got 1.34 x 10^-12 I dont know how to go about the second portion of this question.

Thanks for taking the time to help me!
1
Expert's answer
2012-11-12T11:33:47-0500
you can find escape velocity as from
v^2/r = GM/r^2

v = sqrt(GM/r)
So for red giant we have
v = sqrt ( 6.67*10^(-11)*1.99*10^(30) / (100*7*10^(8)) ) = 6*10^(4) m/s
For planetary nebula we assume mass to be approximately 1M and radius to be
approximately 10^(15) m. So
v = sqrt(6.67*10^(-11)*1.99*10^(30)/(10^(15) ) = 3.3 * 10^2 m/s

You can see that it is much bigger for red giant